Solve the inequality:
2x^2+1<9x-3

Answer by lwsshak3(11628)

Solve the inequality:
2x^2+1<9x-3
2x^2-9x+1+3<0
2x^2-9x+4=0
(x-4)(2x-1)=0
number line:
<...+...1/2...-....4....+....>
solution: (1/2,4)